A dairy is required to have a waste holding pond for runoff. Tom plans to build a pond that will be 20 ft by 15 ft. there will be a border of equal length around the pond. the total area of the pond and the border is 500ft. How wide will the border be?
Answer by mananth(16946)

the pond is 20 by 15
..
let the border width be x
..
so the width including border will be 15+2x ( x on either side)
The length including width = 20+2x
..
Area of entire plot including border = 500 sq.ft= (15+2x)(20+2x)
..
4x^2+70x+300=500
4x^2+70x-200=0
4x^2+80x-10x-200=0
4x(x+20)-10(x+20)=0
(x+20)(4x-10)=0
4x-10=0
4x=10
x=10/4
x= 2.5 feet.
..
CHECK
(15+5)*(20+5)= 500